# BlockFUL: Enabling Unlearning in Blockchained Federated Learning

### v2

@@ VI-E2 Unlearning effectiveness of multi-label unlearning
When #$C_{f}$ (i.e., the number of label categories to be unlearned) is set to 2, 4, and 7 across all models on CIFAR-10, the $AD_{f}$ metric for re-training achieves 0, while the average $AD_{r}$ metric reaches 94.08%. For gradient ascent, the $AD_{f}$ metric stands at 6.64%, and the average $AD_{r}$ metric is 48.77%. On Fashion-MNIST, experimental results show that the $AD_{f}$ metric for re-training attains 0, with an average $AD_{r}$ metric as high as 96.72%; the $AD_{f}$ metric for gradient ascent is 1.64%, and the average $AD_{r}$ metric is 61.1%.
Regarding the $AD_{f}$, the performance of gradient ascent is equally impressive compared to that of re-training the model. Regarding the $AD_{r}$, however, its performance exhibits a downward trend. This suggests that during optimization, the model’s emphasis on unlearned labels leads to insufficient or improper adjustments for retained labels, resulting in decreased performance on $AD_{r}$ due to incomplete learning of relevant knowledge for these labels.
VI-E3 Weight distance analysis
We conduct an analysis of the ResNet18 model using the Fashion-MNIST dataset. In the presence of unlearned data (#$\#C_{f}=1$), we evaluate the distance trend of the unlearning method. Additionally, we compute the layer-wise distances between the re-trained model and the original model, as well as between the gradient ascent method and the original model. We select the top-10 layer-wise examples. As shown in Fig. 5, the gradient ascent model has a similar trend to the re-trained model, proving the effectiveness of the gradient ascent method.
VI-E4 Unlearning speed
Experimental results demonstrated in Tables III and IV reveal that, while sequential unlearning with re-training offers more robust performance on unlearning effectiveness and retained model utility, parallel unlearning with gradient ascent is more efficient. On the CIFAR-10 dataset, as the number of inherited models increases, the cumulative time for unlearning using the gradient ascent method is reduced by a factor of 19, compared to the re-training method. Similarly, on the Fashion-MNIST, the gradient ascent method also demonstrates its efficiency, with an unlearning speed 12 times faster than re-training.
Figure 5: The unlearned data category is #$\#C_{f}=1$. (a) and (b) depict the weight distances for ResNet18 on $r_{5}$ and $r_{8}$, respectively.
VI-F Cost Evaluation
We analyze the time overhead of blockchain execution under the sequential paradigm with re-training and parallel paradigm with gradient ascent for the design goal G4. For each FL task, we establish a blockchain utilizing a single-chain structure. The network utilizes the BFT-based consensus mechanism.
Figure 6: (a) the cost time of multiple start models; (b) the cost time of a single start model.
Figure 7: (a) the cost time of multiple blocks of inheritance models; (b) the cost time of a single block of inheritance model.
Multi-model parallel update cost time. For both re-training and gradient ascent methods, we analyze the blockchain overhead incurred when multiple models are executed in parallel. We assume a DAG graph with only two models and the two models in different blocks, as illustrated in Fig. 6a. It is observed that both the re-training and gradient ascent methods demonstrate linear growth of the time overhead. This arises from the inherent competition for network and computational resources within the BFT-based consensus. The gradient ascent method incurs a maximum reduction of 124% in time overhead, compared to the re-training method.
Chameleon hash and consensus cost time. We conduct multiple trials to measure the CH and consensus computational overhead time for a single model, as illustrated in Fig. 6b. The execution time of CH and consensus remains consistently within a constant range. The consensus time is 92% longer than CH time, contributing to the majority of the blockchain time overhead.
$Tx$-block cost time. When a block contains a single transaction, we conduct experiments across multiple instances across different blocks. The horizontal axis indicates the number of transactions (the number of unlearning models), and the vertical axis gives the time cost. The results are illustrated in Fig. 7a, where we observe a linear growth in time consumption as the number of inheritance models increases for both methods. Compared to re-training, gradient ascent achieves a performance improvement of up to 75.8%.
$Txs$-block cost time. When a block contains multiple transactions, our experiments, aligned with (24) and depicted in Fig. 7b, reveal that the time consumption for both methods increases linearly with the number of inherited models. However, the rate of increase is slower compared to the scenario shown in Fig. 7a, where transactions are packed in different blocks. The improvements in time efficiency are notable, reaching up to 4.4% and 12.6% for the re-training and gradient ascent methods, respectively. This is due to the overhead incurred by multiple block updates. The gradient ascent method achieves an improvement of up to 78.1% compared to re-training. This is attributed to the fact that the parallel paradigm only needs consensus once, whereas the sequential paradigm imposes substantial overhead with every model update from the client, owing to the consensus requirement.
Chameleon hash overhead. We assess the computational overhead of CH on the live chain, modifying the hash values of model $URI$ for nodes $r_{4}$, $r_{5}$, $r_{6}$, and $r_{8}$ as structured in Fig. 4. Each transaction aligns with one block as defined by (19), with $K=N=4$. As shown in Fig. 8a, the modification time for each node is brief and falls within an acceptable range. The computational overhead increases linearly with the number of references, and the costs of both methods are comparable.
Figure 8: (a) represents chameleon hash cost time; (b) represents consensus cost time.
Blockchain consensus overhead. We conduct experiments employing the BFT-based consensus on the DAG structure illustrated in Fig. 4, as demonstrated in Fig. 8b. For the re-training method, our experiments reveal that the time required also rises as the reference number increases. However, for gradient ascent, we only need to conduct consensus once, so the overhead does not increase with the reference count. Notably, in comparison to the CH overhead, the consensus overhead dominates in the majority of the cases. Thus, we have the flexibility to choose different consensus algorithms based on security and efficiency requirements.
VII Conclusion
@@
In response to RQ1 on how Blockchained FL frameworks can unlearn historical models while reducing the costs of updating multiple inherited models, we developed a new framework termed BlockFUL, which employs a dual-chain structure to enable traceable and trustworthy unlearning capability in BLockchained FL. To tackle RQ2 regarding adaptable unlearning methods for Blockchained FL, we presented two paradigms under the BlockFUL framework: parallel and sequential unlearning, utilizing the prevalent unlearning methods of gradient ascent and re-training for demonstration. Our analysis revealed that parallel unlearning with gradient ascent is more cost-effective than sequential unlearning, particularly in consensus and transmission costs, as the number of model updates increases. Our experimental results demonstrated that sequential unlearning with re-training maintains high accuracy for retained data, while the effectiveness of parallel unlearning varies with inheritance depth and model differences.

### v3

@@ VII-E2 Unlearning effectiveness of multi-label unlearning
When $C_{f}$ (i.e., the number of label categories to be unlearned) is set to 2, 4, and 7 across all models on CIFAR-10, the $AD_{f}$ metric for gradient ascent is 6.64%, and the average $AD_{r}$ is 48.77%. On Fashion-MNIST, the experimental results show that the $AD_{f}$ for gradient ascent is 1.64%, and the average $AD_{r}$ is 61.1%. Gradient ascent performs similarly to re-training on $AD_{f}$, but its declining $AD_{r}$ suggests that focusing on unlearning labels hampers retention of other labels, affecting overall performance.
Different models and datasets balance computational cost and accuracy differently. CIFAR-10, with its larger dataset, demands more re-training resources than Fashion-MNIST but achieves higher accuracy. Model complexity also matters—ResNet18, being more complex than MobileNetV2, requires higher re-training costs but delivers superior accuracy. In gradient ascent, ResNet18 maintains its accuracy advantage over MobileNetV2 despite similar computational costs, highlighting the benefits of model complexity in this method. Thus, model and dataset selection should match task requirements: MobileNetV2 suits resource-constrained scenarios (e.g., real-time applications), while ResNet18 is ideal for accuracy-critical tasks (e.g., image recognition).
Takeaway—balancing unlearning accuracy and scalability. Re-training offers perfect forgetting with high retention, particularly for lightweight models like AlexNet. Gradient ascent, while less effective in retention, enables cost-efficient unlearning with acceptable forgetting performance, and benefits from model complexity. These results support flexible method selection based on task constraints such as accuracy demand and computational resources.
VII-E3 Non-IID and IID
Tables III and IV present experimental results under IID and non-IID data distributions, where client data is either sample-non-IID or label-non-IID. In the sample-non-IID scenario, clients share the same classes but with different data within each class; in the label-non-IID scenario, clients have disjoint class distributions (e.g., client 1 does not have classes 70-74, and client 2 does not have classes 60-64). For IID data, clients 1 and 2 share both labels and data. It is observed that the GA method effectively achieves unlearning in most cases, with particularly better performance under the sample-non-IID setting. As $C_{f}$ increases, GA accuracy under IID data drops significantly because sample-non-IID maintains greater data diversity, enhancing model performance, whereas label-non-IID’s disjoint class distribution can negatively impact learning.
Tables V and VI present the experimental results of non-IID unlearning performance under different label settings on CIFAR-100 (ResNet18) and Yahoo! Answers (Bert-Base-Cased), respectively. We observe that as the number of label-non-IID labels increases (e.g., in the CIFAR-100 dataset (ResNet18), where the number of inconsistent labels increases from 5 to 15), the unlearning performance remains relatively high across different models and datasets.
Beyond re-training and gradient ascent, we explored distillation and fine-tuning. As shown in Tables VII and VIII, distillation slightly underperforms re-training in accuracy while incurring higher time overhead, and fine-tuning proves ineffective. Fine-tuning updates only the final layers of a pre-trained model on a dataset excluding the unlearned data. To improve unlearning and minimize $AD_{f}$, we applied a large learning rate, effectively removing unlearned data but causing instability due to distribution shifts, making it harder to retain essential information [31].
TABLE VII: Fine-tuning and Distill Unlearning Performance on CIFAR-100
Model $C_{f}$ Metrics Original (%) Fine-tuning (%) Distill (%) Cumulative Unlearning Time (s)
Fine-tuning Distill
$r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$
AlexNet
1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 1.03 1.03 3.21 81.74 86.25 86.76 0.85 1.16 2.45 83.14 154.61 296.74
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
10 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 1.13 4.77 67.04 80.44 81.07 91.32 0.84 2.14 3.31 76.31 149.60 329.95
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 52.73 0.00 0.00 0.00
ResNet18
1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 1.03 2.44 1.03 84.93 84.93 81.32 1.04 2.32 3.74 71.31 131.44 250.61
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
10 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 6.87 1.68 2.04 81.02 79.76 91.04 1.85 2.54 3.84 76.12 152.33 270.34
$A_{D_{f}}\downarrow$ 99.98 99.99 99.99 2.61 0.00 0.00 0.00 0.00 0.00
TABLE VIII: Re-training and Gradient Ascent Unlearning Performance on CIFAR-100
Model $C_{f}$ Metrics Original (%) Re-training (%) Gradient Ascent (%) Cumulative Unlearning Time (s)
Re-training Gradient Ascent
$r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$
AlexNet 1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.13 92.69 96.32 30.14 59.53 91.31 1.34 1.85 2.42
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 1.38 0.00 0.00
10 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 43.61 39.30 49.94 29.53 57.16 89.17 1.67 2.14 2.75
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 21.05 17.76 29.31
ResNet18 1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 92.81 79.55 87.06 31.64 62.51 92.64 1.84 2.35 2.94
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 1.38 0.00 0.00
10 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 81.02 79.76 91.04 29.95 59.14 88.18 1.93 2.45 3.01
$A_{D_{f}}\downarrow$ 99.98 99.99 99.99 0.00 0.00 0.00 0.00 0.00 0.00
TABLE IX: Large-scale Federated Unlearning Performance of Language and Graph Models on Yahoo! Answers
Model $C_{f}$ Metrics Original (%) Re-training (%) Gradient Ascent (%) Cumulative Unlearning Time (s)
Re-training Gradient Ascent
$r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$
Bert-Base-Cased 1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 84.40 96.39 84.49 105.63 215.14 313.75 1.97 3.05 4.14
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 3.96 8.91 6.25
2 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 71.48 43.74 30.87 95.61 190.38 308.74 2.07 3.51 4.81
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 15.13 19.45 14.51
4 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 70.47 49.31 50.85 72.16 147.47 232.35 2.57 4.34 5.42
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 12.29 13.63 18.18
GCNN 1 $A_{D_{r}}\uparrow$ 99.56 99.93 96.43 99.69 99.99 97.64 58.42 44.06 42.48 35.14 71.36 107.25 1.02 2.02 3.17
$A_{D_{f}}\downarrow$ 98.81 99.99 94.64 0.00 0.00 0.00 14.79 8.33 20.71
2 $A_{D_{r}}\uparrow$ 99.56 99.93 96.43 99.99 99.99 98.31 60.57 41.48 40.27 30.66 61.24 92.34 1.44 2.87 3.86
$A_{D_{f}}\downarrow$ 99.11 99.99 95.60 0.00 0.00 0.00 17.80 8.01 13.19
4 $A_{D_{r}}\uparrow$ 99.56 99.93 96.43 99.99 99.99 98.34 54.34 35.80 41.72 23.31 46.46 71.87 2.04 3.14 4.51
$A_{D_{f}}\downarrow$ 99.07 99.99 95.67 0.00 0.00 0.00 10.54 10.00 11.19
TABLE X: Federated Unlearning Performance on TinyImageNet
Model $C_{f}$ Metrics Original (%) Re-training (%) Gradient Ascent (%) Cumulative Unlearning Time (s)
Re-training Gradient Ascent
$r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$ $r_{5}$ $r_{6}$ $r_{8}$
ResNet18 1 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 99.98 99.98 99.99 114.95 220.39 320.29 21.44 25.78 28.62
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 0.00 1.44 18.60
4 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 83.25 93.73 81.06 111.89 213.31 316.42 20.67 24.97 27.51
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 7.16 21.49 14.33
6 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.96 99.99 99.99 78.36 59.99 68.41 109.39 212.12 311.32 20.13 24.56 26.47
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 12.94 11.27 19.34
10 $A_{D_{r}}\uparrow$ 99.99 99.99 99.99 99.99 99.99 99.99 57.39 65.26 56.41 106.21 205.36 308.74 18.30 21.67 23.68
$A_{D_{f}}\downarrow$ 99.99 99.99 99.99 0.00 0.00 0.00 12.81 13.30 12.80
Table IX presents multi-label unlearning results on the Yahoo! Answers dataset using the Bert-Base-Cased and GCNN models with 50 clients. As client numbers increase, $AD_{f}$ slightly rises (11.19% to 20.71%) but remains low overall. BlockFUL achieves better time efficiency in GA than re-training, despite a minor trade-off in $AD_{r}$, with GA significantly reducing computational costs. Similar trends appear in Table X for TinyImageNet, where GA consistently outperforms re-training. The $AD_{r}$ values remain within an acceptable range, ensuring the method’s effectiveness.
Takeaway—scalable unlearning under non-IID and federated settings. Gradient ascent maintains strong unlearning performance across IID and non-IID scenarios, particularly excelling under sample- non-IID due to preserved data diversity. It remains effective as the number of forgotten labels increases and scales well across federated setups, offering significant time savings over re-training with only minor trade-offs in retention accuracy.
VII-F Evaluating Unlearning Scalability
Figure 5: (a) and (b) show the sequential throughput and latency; (c) and (d) show the parallel throughput and latency.
Figure 6: (a) and (b) show the sequential throughput and latency across different DAG depths; (c) and (d) show the parallel throughput and latency across different DAG depths.
Figure 7: (a) and (b) show the throughput and latency of sequential unlearning across different datasets and models; (c) and (d) show the throughput and latency of parallel unlearning across different datasets and models.
We analyze FL and blockchain time overhead under sequential (re-training) and parallel (gradient ascent) paradigms for G4. Each unlearning request maps to a model chain of depth 1 to 8, with each model linked to a block. We evaluate performance under sequential (100–800 requests) and parallel (1000–8000 requests) paradigms across varying scales and processing strategies.
In the sequential paradigm, throughput decreases as the number of unlearning requests grows, with both PBFT and PoW consensus mechanisms affected (Fig. 5). PoW experiences a steeper decline beyond 500 requests due to its higher block overhead, while PBFT sustains more stable throughput thanks to its lightweight consensus process. Latency, however, rises with increasing requests (Fig. 5), with PoW exhibiting the highest latency due to its complex operations, and PBFT with block requests achieving the lowest, reflecting its efficiency in handling sequential tasks.
The parallel paradigm significantly improves scalability, with throughput increasing as unlearning requests grow, due to concurrent handling of FL and block tasks (Fig. 5). PBFT outperforms PoW beyond 5000 requests by better managing block overhead, while PoW’s heavier consensus limits its performance. Latency drops with more requests (Fig. 5) as parallelism reduces processing time per request, though PoW still shows the highest latency compared to PBFT.
Moreover, DAG depth and model/dataset complexity have a non-neglible impact. In the sequential paradigm, throughput decreases with depth—PoW drops sharply beyond depth 5—while latency peaks at depth 8 (Figs. 66). In contrast, parallel processing yields higher throughput and lower latency, with PBFT consistently outperforming PoW, especially at depth 8 (Figs. 66). For different models and datasets, sequential throughput drops in complex cases like ResNet18 on TinyImageNet beyond 500 requests, with increased latency (Figs. 77), while parallel processing maintains lower latency and higher throughput, particularly for simpler setups like AlexNet on CIFAR100 after 5000 requests (Figs. 77).
Takeaway—parallel paradigm scales efficiently with reduced latency. Parallel unlearning sig- nificantly improves throughput and lowers latency under high request volume and deep model inheritance, especially with high-speed consensus (e.g., PBFT). Unlike sequential re-training, which suffers from increased overhead and latency as DAG depth and request count grow, gradient ascent with parallel execution maintains stable performance and scalability across models, datasets, and blockchain protocols.
VIII Conclusion And Further Work
@@
In response to RQ1 on minimizing update costs for unlearning historical models in Blockchained FL, we developed BlockFUL, a dual-chain framework ensuring traceable and trustworthy unlearning. For RQ2 on adaptable unlearning methods, we introduced two paradigms in BlockFUL: parallel unlearning with gradient ascent and sequential unlearning with re-training. Our analysis showed that parallel unlearning is more cost-effective, particularly for consensus and transmission costs as model updates grow. Experiments revealed that sequential unlearning maintains high accuracy for retained data, while parallel unlearning effectiveness depends on inheritance depth and model variations. In future work, we plan to explore more efficient algorithms, such as model compression, to reduce computational costs for unlearning in large-scale datasets and complex models without compromising accuracy.